# Reduce-Rank Matrix Integer-Valued Autoregressive Model

### v1

@@ 1 Introduction
where $\{e_{t}\}$ is an i.i.d. Bernoulli random sequence with the same parameter $\lambda$. Alzaid and Al-Osh (1988) and Al-Osh and Aly (1992) developed INAR(1) models whose marginal distributions follow the geometric, generalized Poisson, and negative binomial distributions, respectively. Jin-Guan and Yuan (1991) extended the INAR(1) model to a higher-order INAR(p) process and investigated its robustness and parameter estimation methods. Bu (2008) studied the estimation problem for the INAR(p) model and proposed the maximum likelihood estimator. The study compared the performance of this estimator with the conditional least squares (CLS) and conditional maximum likelihood (CML) estimators for the Poisson INAR(2) model. To better analyze real-world data, the INAR model has been expanded in more directions. For example, Jazi et al. (2012) proposed the first-order zero-inflation integer autoregressive (ZI-INAR(1)) model to process zero-inflation data, and the experimental effect on the Pittsburgh drug crime data was good. Pang et al. (2023) proposed a time-varying parameter integer autoregressive (TV-INAR) model, considering the parameter $\alpha$ varying over time, and showed its improved fit over conventional INAR models when applied to sexual offense data in Barina City, Australia.
@@
The INAR model is unable to adapt to the diversified development of data; therefore Franke and Rao Subba (1993) introduced a significant multivariate INAR(1) model constructed using a matrix of univariate independent binomial thinning operators. Subsequently, Latour (1997) proposed another important extension: the multivariate generalized integer autoregressive (MGINAR) model. This model replaces the binomial thinning operators with generalized thinning operators based on counting random variables that are not necessarily Bernoulli-distributed. Kirchner (2017) further extended the multivariable INAR(1) model to the multivariable INAR(p) model, using CLS to estimate the parameters, and proved the consistency and asymptotic normality of the estimators. Pedeli and Karlis (2011) and Pedeli and Karlis (2013) successively studied the properties and estimation problems of multivariate INAR models where the error terms follow bivariate and multivariate Poisson distributions, respectively. To address over-dispersion in multivariate integer-valued data, Yu and Zheng (2024) proposed a multivariate INAR(1) model with a mixed innovation term, featuring multivariate log-normal and multivariate log-Poisson distributions. To effectively model multivariate integer-valued time series with periodic characteristics, Santos et al. (2021) proposed a periodic multiple integer autoregressive (PMINAR(1)) model, incorporating periodic time-varying parameters and multiple negative binomial innovations. These studies have further promoted the development of integer-valued time series analysis.
@@
Due to the increasing complexity and dynamic nature of data, the scope of analysis has broadened to incorporate multiple variables observed across various entities. These data can be structured in matrix form, as exemplified by economic indicators such as the gross domestic product (GDP) of multiple countries, bank interest rates, and total industrial production.Over time, the continuous observation of such matrix-form data gives rise to a matrix time series. This concept was first elaborated in the signal processing literature by (Walden and Serroukh (2002)). In recent years, the academic community has demonstrated growing interest in the modeling and analysis of this type of time series and has achieved a series of innovative results.
@@
The modeling and analysis of matrix time series can be broadly categorized into two approaches. The first category comprises methods based on vectorizing the matrix data (Basu and Michailidis (2015), Bauer and Vorkink (2011)). These methods adapt the data to the vector autoregressive (VAR) model by restructuring it, yet this process results in a loss of structural information and introduces an excessive number of model parameters. The second category of methods preserves the original data structure. These approaches utilize matrix factorization techniques to construct models directly from the matrix observations (Wang et al. (2019), Chen et al. (2020)). The first-order matrix autoregressive (MAR(1)) model proposed by Chen et al. (2021) deeply describes the dynamic interaction between rows and columns in the matrix time series. Compared with VAR model ((Hannan (1970), Lütkepohl (2005)), MAR method significantly reduces the number of the model parameters, and enhances the structural information. Han et al. (2024) further proposed a model based on the bilinear transformation, which simplifies the model structure. All the above studies are conducted within the real number domain. Even so, the above-mentioned progresses have opened up a new path for the theoretical and application research of matrix integer-valued time series. Recently, Xu (2024) introduced a matrix integer-valued autoregressive (MINAR) model and employed both the projection estimation method and the iterative conditional least squares estimation (ICLSE) method for parameter estimation.
@@
As data dimensions increase, redundant information impairs modeling performance and estimation efficiency, making it necessary to employ alternative methods to mitigate the curse of dimensionality. Although the matrix sequence can be stacked column by column and transformed into a vector sequence for dimensionality reduction or thinning processing, e.g. the regularization and thinning regression techniques(Basu and Michailidis (2015) and Kock2015), and dimensionality reduction techniques, such as the reduce-rank vector autoregressive (RRVAR) model (Cubadda and Hecq (2021)), blind stacking will destroy the matrix data structure and lose the mutual influence of specific dimensions. To reduce the number of parameters without destroying the data structure, Xiao et al. (2022), Liu et al. (2023), and others proposed a reduced-rank matrix autoregressive (RRMAR) model based on the MAR model. They applied rank reduction to its bilinear coefficient matrix to mitigate the effects of high dimensionality and noise. Regrettably, the above mentioned methods require further research in the field of integer time series, which limits the application of integer time series in the high-dimensional integer field.
@@ 5.3 Crime data analysis
Based on the selected ranks and the observation matrices in the training dataset, the MGINAR(1), MINAR(1) and RRMINAR(1) models are established for one-step-prediction, respectively. The training and prediction performance is evaluated using the following four error metrics:
@@
Note that, with the exception of MINAR(1) and RRMINAR(1), the parameter matrix estimates of the remaining models may contain a small number of elements that are slightly negative. This makes the relevant models unable to produce predictions and renders them practically meaningless. However, since the negative values are small, it can be considered that the corresponding components contribute only minimally to the model predictions. These negative values can be corrected by taking their absolute values or setting them to zero, thereby obtaining parameter matrix estimates with entirely positive values. Here, we take their absolute values.

### v2

@@ 1 Introduction
where $\{e_{t}\}$ is an i.i.d. Bernoulli random sequence with the same parameter $\lambda$. Alzaid and Al-Osh (1988) and Al-Osh and Aly (1992) developed INAR(1) models whose marginal distributions follow the geometric, generalized Poisson, and negative binomial distributions, respectively. Jin-Guan and Yuan (1991) extended the INAR(1) model to a higher-order INAR(p) process and investigated its robustness and parameter estimation methods. Bu and McCabe (2008) studied the estimation problem for the INAR(p) model and proposed the maximum likelihood estimator. The study compared the performance of this estimator with the conditional least squares (CLS) and conditional maximum likelihood (CML) estimators for the Poisson INAR(2) model. To better analyze real-world data, the INAR model has been expanded in more directions. For example, Jazi et al. (2012) proposed the first-order zero-inflation integer autoregressive (ZI-INAR(1)) model to process zero-inflation data, and the experimental effect on the Pittsburgh drug crime data was good. Pang et al. (2023) proposed a time-varying parameter integer-valued autoregressive (TV-INAR) model that allows the parameter $\alpha$ to vary over time. The model demonstrated a better fit than conventional INAR models when applied to sexual offense data from Barina City, Australia.
@@
The INAR model is unable to adapt to the diversified development of data; therefore Franke and Rao Subba (1993) introduced a significant multivariate INAR(1) model constructed using a matrix of univariate independent binomial thinning operators. Subsequently, Latour (1997) proposed another important extension: the multivariate generalized integer autoregressive (MGINAR) model. This model replaces the binomial thinning operators with generalized thinning operators based on counting random variables that are not necessarily Bernoulli-distributed. Kirchner (2017) further extended the multivariate INAR(1) to a $p$-th order model, estimating parameters via CLS and establishing the consistency and asymptotic normality of the estimators. Pedeli and Karlis (2011) and Pedeli and Karlis (2013) investigated the properties and estimation of multivariate INAR models with bivariate and multivariate Poisson-distributed error terms, respectively. To address over-dispersion in multivariate integer-valued data, Yu and Zheng (2024) proposed a multivariate INAR(1) model with a mixed innovation term, featuring multivariate log-normal and multivariate log-Poisson distributions. To effectively model multivariate integer-valued time series with periodic characteristics, Santos et al. (2021) proposed a periodic multiple integer autoregressive (PMINAR(1)) model, incorporating periodic time-varying parameters and multiple negative binomial innovations. These studies have further promoted the development of integer-valued time series analysis.
@@
Due to the increasing complexity and dynamic nature of data, the scope of analysis has broadened to incorporate multiple variables observed across various entities. These data can be structured in matrix form, as exemplified by economic indicators such as the gross domestic product (GDP) of multiple countries, bank interest rates, and total industrial production. Over time, the continuous observation of such matrix-form data gives rise to a matrix time series. This concept was first elaborated in the signal processing literature by Walden and Serroukh (2002). In recent years, the academic community has demonstrated growing interest in the modeling and analysis of this type of time series and has achieved a series of innovative results.
@@
The modeling and analysis of matrix time series can be broadly categorized into two approaches. The first category comprises methods based on vectorizing the matrix data, such as Basu and Michailidis (2015) and Bauer and Vorkink (2011). These methods reshape the data to fit a vector autoregressive (VAR) model, but this leads to both a loss of structural information and over-parameterization. The second category of methods preserves the original data structure. These approaches utilize matrix factorization techniques to construct models directly from the matrix observations (Wang et al. (2019), Chen et al. (2020)). The first-order matrix autoregressive (MAR(1)) model proposed by Chen et al. (2021) deeply describes the dynamic interaction between rows and columns in the matrix time series. Compared with VAR model ((Hannan (1970), Lütkepohl (2005)), MAR method significantly reduces the number of the model parameters, and enhances the structural information. Han et al. (2024) further proposed a model based on the bilinear transformation, which simplifies the model structure. All the above studies are conducted within the real number domain. Even so, the above-mentioned progresses have opened up a new path for the theoretical and application research of matrix integer-valued time series. Recently, Xu (2024) introduced a matrix integer-valued autoregressive (MINAR) model and employed both the projection estimation method and the iterative conditional least squares estimation (ICLSE) method for parameter estimation.
@@
As data dimensions increase, redundant information impairs modeling performance and estimation efficiency, making it necessary to employ alternative methods to mitigate the curse of dimensionality. Although the matrix sequence can be stacked column by column and transformed into a vector sequence for dimensionality reduction or thinning processing, e.g. the regularization and thinning regression techniques(Basu and Michailidis (2015) and Bauer and Vorkink (2011)), and dimensionality reduction techniques, such as the reduce-rank vector autoregressive (RRVAR) model (Cubadda and Hecq (2021)), blind stacking will destroy the matrix data structure and lose the mutual influence of specific dimensions. To reduce the number of parameters without destroying the data structure, Xiao et al. (2022), Liu et al. (2023), and others proposed a reduced-rank matrix autoregressive (RRMAR) model based on the MAR model. They applied rank reduction to its bilinear coefficient matrix to mitigate the effects of high dimensionality and noise. Regrettably, existing methods in integer-valued time series remain underdeveloped, restricting their applicability in high-dimensional contexts.
@@ 5.3 Crime data analysis
Based on the selected ranks and the observation matrices in the training dataset, the MGINAR(1), MINAR(1) and RRMINAR(1) models are established for one-step-prediction, respectively. The training and prediction performance is evaluated using the following four error metrics, with any zero denominator in $\rm E_{3}$ or $\rm E_{4}$ replaced by a value of 1:
@@
Note that, except for MINAR(1) and RRMINAR(1), the parameter estimates of other models may contain a small number of slightly negative elements. This makes the relevant models unable to produce predictions and renders them practically meaningless. However, the small magnitude of these negative values implies that their contribution to the model’s predictions is marginal. The negative values can be corrected by taking absolute values or setting them to zero to obtain fully non-negative parameter estimates. Here, we take their absolute values.